Kendra is 6 years older than one half Gitta's age. Their age sum is 57. How old are Kendra and Gitta?
Answer by macston(5194)

.
G=Gitta's age; K=Kendra's age=0.5G+6yrs
G+K=57yrs
G+0.5G+6yrs=57yrs
1.5G=51yrs
G=34
ANSWER 1: Gitta is 34 years old.
.
K=0.5G+6yrs=0.5(34yrs)+6yrs=17yrs+6yrs=23yrs
ANSWER 2: Kendra is 23 years old.
.
CHECK:
G+K=57 yrs
34 yrs +23 yrs = 57 yrs
57 yrs = 57 yrs